PT AKR company purchased a machine on january 1, 2011 for Rp 3,000,000. At the date acquisition, the machine had a life of six years with no residual value. The machine is being depreciated on a straight line basis. On January 1 2014, AKR determined that the machine had a useful life of eight years from date of acqusition with no residual value. What should be the depreciation for the year 2014?
A. 300,000
B. 500,000
C. 187,500
D. 250,000
Answer #1

Solution:

Depreciation charged from 2011 to 2013 = 3,000,000 * 3/6 = 1,500,000

Carrying value of machine on 01.01.2014 = 3,000,000 - 1,500,000 = 1,500,000

Remaining useful life of machine = 8-3 = 5 years

Depreciation for the year 2014 = 1,500,000 / 5= 300,000

Hence option A is correct,